### v1

@@ 4 Numerical Experiments
To evaluate the proposed method comprehensively, it is insufficient to solely rely on a single indicator. Before providing the assessment criteria, it is necessary to establish the confusion matrix for the binary classifier. The confusion matrix provides fundamental details on the classifier’s actual and expected classifications (see Table 1). The discriminant criteria we offer include accuracy, precision, recall, sensitivity, specificity, and F1 score.
@@ 4.1 Experiments on Different Datasets
In this section, to illustrate the universality of the proposed framework, we compare our proposed generalized framework, which combines standard first-order optimization algorithms SGD and second-order optimization algorithms oBFGS and oNAQ, with these algorithms on 12 real datasets. These datasets are diverse, including sparse, dense, imbalanced, high-dimensional, and low-dimensional. They were collected from the University of California, Irvine (UCI) Machine Learning repository [^1] https://archive.ics.uci.edu, Libsvm data [^2] https://www.csie.ntu.edu.tw/~cjlin/libsvmtools/datasets/. Details of the features are shown in Table 2. The second column indicates the source of the data set, the third column indicates the quantity of data, the fourth column indicates the dimension, and the fifth column indicates the quantity of data categories. The proposed approach differs from the standard approach only in whether the weights are set and updated.
@@
The variance $\sigma$ in the AW function is set to 1 empirically. We conducted 100 iterative experiments, taking the total amount of iterations as the horizontal coordinate and the test accuracy as the vertical coordinate to depict the performance of each algorithm, as shown in Fig.6. Fig.5(a)-5(c) respectively plotted the experimental results of the 12 datasets in Table 2. No matter whether the proposed method in the initial result performs well or not, it tends to be stable after a certain number of iterations. We can conclude that the proposed approach effectively improves the performance of SGD, oBFGS and oNAQ on all 12 datasets.
We additionally evaluate the efficacy of the proposed framework, AW-WSVM, by using the confusion matrix. To ensure the stability of the results, we iterated 100 times. The metrics of different algorithms obtained from the confusion matrix can be presented in Table 4-8, with the highest metrics bolded. Conclusions are easy to come by, and the proposed method outperforms SGD, oBFGS, and oNAQ on 11 datasets. The proposed method exhibits lower results with regard to accuracy, recall, sensitivity, and F1 scores when compared to oBFGS without weight involvement on the A7a dataset. However, when considering the entirety of the recommended technique, it has successfully attained the intended outcome.
@@
The findings presented in Table 4-8 indicate that, when considering a specific algorithm applied to the identical dataset, the recall rate, precision, sensitivity, and F1 score remain consistent. However, the prediction accuracy demonstrates improvement. Our model demonstrates a higher level of accuracy in identifying samples from each category throughout data processing. Put simply, the model exhibits enhanced generalization and stability, enabling it to effectively process data in real-world scenarios.
4.2 Statistical Comparison
$\chi_{F}^{2}=\frac{12N}{K(K+1)}[R_{j}^{2}-\frac{(K(K+1)^{2})}{4}]$ (8)
where $K$ is the number of algorithms involved in the comparison.
Table 5 shows the average ranking $R_{j}$ of the six classifiers in test accuracy. The p-value of Friedman’s test is $1.7122\times 10^{-9}$. The P-value is less than the threshold for significance level $\alpha=0.05$, indicating that the null assumption should be discarded. The above results show that there are great differences among all classification methods from a statistical point of view. We compared the accuracy of the six classifiers (see Table 9), with a p-value of $1.8302\times 10^{-8}$ for Friedman’s test, indicating significant differences among all classification methods.
@@ 4.2 Statistical Comparison
Fig.6(a) and Fig.6(b) respectively report the analysis results of six classifiers in terms of test accuracy and precision. The test accuracy of the standard SGD, oBFGS, and oNAQ algorithms does not exhibit any substantial variation, but there is a notable disparity between the proposed approach in this research. Furthermore, our analysis indicates that the integration of all three algorithms with AW-WSVM leads to enhanced algorithm performance. This improvement is evident from the overall higher ranking of the combined algorithms across all datasets, showcasing varied progress. The results indicate a negligible disparity in the classification accuracy across standard SGD, oBFGS, and oNAQ. Each of the three algorithms was examined independently. The performance of SGD was enhanced by using the weight update approach suggested in this research (AW-WSVM+SGD). Notably, there were notable disparities between the two algorithms when compared to the original SGD. The same applies to second-order approaches such as BFGS and NAQ.
Table 5: Precision of six methods on different datasets.
Datasets oNAQ AW-WSVM+oNAQ oBFGS AW-WSVM+oBFGS SGD AW-WSVM+SGD
A7a 0.8201 0.8216 0.8212 0.8200 0.7667 0.8486
A8a 0.8250 0.8250 0.8219 0.8275 0.7735 0.8282
A9a 0.8206 0.8216 0.8181 0.8222 0.7710 0.8258
Mushroom 0.9651 0.9798 0.9564 0.9714 0.7585 0.9725
Yeast 0.6241 0.6285 0.6277 0.6293 0.6004 0.6004
Ijcnn1 0.9050 0.9180 0.9050 0.9169 0.9050 0.9133
W1a 0.9763 0.9802 0.9747 0.9787 0.9706 0.9777
W2a 0.9750 0.9802 0.9733 0.9813 0.9704 0.9813
W3a 0.9746 0.9788 0.9744 0.9809 0.9702 0.9820
W4a 0.9730 0.9809 0.9722 0.9815 0.9702 0.9824
W5a 0.9729 0.9827 0.9749 0.9813 0.9699 0.9825
W6a 0.9753 0.9827 0.9732 0.9847 0.9707 0.9842
Table 6: Recall of six methods on different datasets.
Datasets oNAQ AW-WSVM+oNAQ oBFGS AW-WSVM+oBFGS SGD AW-WSVM+SGD
A7a 0.8201 0.8216 0.8212 0.8200 0.7667 0.8486
A8a 0.8250 0.8250 0.8219 0.8275 0.7735 0.8282
A9a 0.8206 0.8216 0.8181 0.8222 0.7710 0.8258
Mushroom 0.9651 0.9798 0.9564 0.9714 0.7585 0.9725
Yeast 0.6241 0.6285 0.6277 0.6293 0.6004 0.6004
Ijcnn1 0.9050 0.9180 0.9050 0.9169 0.9050 0.9133
W1a 0.9763 0.9802 0.9747 0.9787 0.9706 0.9777
W2a 0.9750 0.9802 0.9733 0.9813 0.9704 0.9813
W3a 0.9746 0.9788 0.9744 0.9809 0.9702 0.9820
W4a 0.9730 0.9809 0.9722 0.9815 0.9702 0.9824
W5a 0.9729 0.9827 0.9749 0.9813 0.9699 0.9825
W6a 0.9753 0.9827 0.9732 0.9847 0.9707 0.9842
Table 7: Sensitivity of six methods on different datasets.
Datasets oNAQ AW-WSVM+oNAQ oBFGS AW-WSVM+oBFGS SGD AW-WSVM+SGD
A7a 0.8201 0.8216 0.8212 0.8200 0.7667 0.8486
A8a 0.8250 0.8250 0.8219 0.8275 0.7735 0.8282
A9a 0.8206 0.8216 0.8181 0.8222 0.7710 0.8258
Mushroom 0.9651 0.9798 0.9564 0.9714 0.7585 0.9725
Yeast 0.6241 0.6285 0.6277 0.6293 0.6004 0.6004
Ijcnn1 0.9050 0.9180 0.9050 0.9169 0.9050 0.9133
W1a 0.9763 0.9802 0.9747 0.9787 0.9706 0.9777
W2a 0.9750 0.9802 0.9733 0.9813 0.9704 0.9813
W3a 0.9746 0.9788 0.9744 0.9809 0.9702 0.9820
W4a 0.9730 0.9809 0.9722 0.9815 0.9702 0.9824
W5a 0.9729 0.9827 0.9749 0.9813 0.9699 0.9825
W6a 0.9753 0.9827 0.9732 0.9847 0.9707 0.9842
Table 8: F1-score of six methods on different datasets.
Datasets oNAQ AW-WSVM+oNAQ oBFGS AW-WSVM+oBFGS SGD AW-WSVM+SGD
A7a 0.8201 0.8216 0.8212 0.8200 0.7667 0.8486
A8a 0.8250 0.8250 0.8219 0.8275 0.7735 0.8282
A9a 0.8206 0.8216 0.8181 0.8222 0.7710 0.8258
Mushroom 0.9651 0.9798 0.9564 0.9714 0.7585 0.9725
Yeast 0.6241 0.6285 0.6277 0.6293 0.6004 0.6004
Ijcnn1 0.9050 0.9180 0.9050 0.9169 0.9050 0.9133
W1a 0.9763 0.9802 0.9747 0.9787 0.9706 0.9777
W2a 0.9750 0.9802 0.9733 0.9813 0.9704 0.9813
W3a 0.9746 0.9788 0.9744 0.9809 0.9702 0.9820
W4a 0.9730 0.9809 0.9722 0.9815 0.9702 0.9824
W5a 0.9729 0.9827 0.9749 0.9813 0.9699 0.9825
W6a 0.9753 0.9827 0.9732 0.9847 0.9707 0.9842
((a)) A7a
((b)) A8a
((c)) A9a
((d)) Mushroom
((e)) Yeast
((f)) Ijcnn1
((g)) W1a
((h)) W2a
((i)) W3a
((j)) W4a
((k)) W5a
((l)) W6a
@@
Table 10: Friedman ranking of these methods according to precisions on different datasets.
@@
A7a 4 2 3 5 6 1
A8a 3.5 3.5 4 2 5 1
A9a 4 2 5 1 6 3
Mushroom 4 1 5 3 6 2
@@
Ijcnn1 4.5 1 4.5 2 4.5 3
W1a 4 1 5 2 6 3
W2a 3 2 4 1.5 5 1.5
W3a 4 3 5 2 6 1
W4a 4 3 5 2 6 1
W5a 5 1 4 3 6 2
W6a 4 3 5 1 6 2
R 4 2.04 4.38 2.13 5.67 2.17
((a)) Classification accuracy
((b)) Classification precision
@@ 4.3 Experiments on Emotional Classification with Different Imbalance Ratio(IR)
In this section, we collect the dataset of emotion classification, which comes from gitte [^3] https://gitee.com/zhangdddong/toutiao-text-classfication-dataset. (see Tabel 11). The last column in the table indicates the imbalance ratio (IR). In daily life, different kinds of news data are always imbalanced, for example, entertainment news will account for more than science and technology news.
@@
In order to further illustrate the effectiveness of the proposed AW-WSVM framework, we process the same data and get G-means under different IRs[43]. In order to further illustrate the effectiveness of the proposed AW-WSVM framework, we process the same data and get G-means under different IR. Taking 102vs100 as an example, we plot the results in Fig8. It is not difficult to find that our proposed method performs best in both low IR and high IR.
@@
5 Conclusion
Considering massive amounts of data, the measurement of the distance between the samples and the decision hyperplane can be calculated with a Gaussian distribution. Therefore, the weights of the samples can be updated by the Adaptive weight function (AW function). Our analysis demonstrates that the weights generated by the AW function obey a probability distribution in each class. Unconstrained soft interval support vector machines have been improved through the addition of weighting coefficients, aligning with the fundamental concept. Subsequently, we can allocate greater weights to samples that are in close proximity to the decision hyperplane, while assigning lower weights to the other samples. Our allocation technique is appropriate since the support vectors, which are placed near the decision hyperplane, correspond with the concept that samples close to the decision plane have higher influence.
The proposed framework, AW-WSVM, was evaluated on a total of 12 datasets sourced from UCI repositories, LIBSVM repositories, and emotional datasets. The experiments conducted confirm that the proposed framework AW-WSVM surpasses the conventional first-order approaches SGD, second-order methods oBFGS, and oNAQ. Furthermore, it exhibited superior performance in terms of classification accuracy, precision, recall, sensitivity, and F1 score. In addition, we also conducted statistical experiments on these six classifiers. Friedman tests and Nemenyi post hoc tests demonstrate that the standard algorithms SGD, oBFGS, and oNAQ all perform better within our proposed framework. We not only do experiments on standard datasets, but also apply the framework to emotion classification datasets. From the G-means and other indicators, the framework proposed in this paper can effectively improve the performance of standard optimization algorithms, no matter whether the imbalance index is high or low. Moreover, the core of the proposed generalized framework, AW-WSVM, is to look for data samples near the classified hyperplane. Consequently, it may be effectively integrated with any soft-margin SVM solution techniques.

### v3

@@ 4 Numerical Experiments
To evaluate the proposed method comprehensively, it is insufficient to solely rely on a single indicator. Before providing the assessment criteria, it is necessary to establish the confusion matrix for the binary classifier. The confusion matrix provides fundamental details on the classifier’s actual and expected classifications (see Table 1). The discriminant criteria we offer include accuracy, precision, recall, specificity, F1 score and G-mean.
@@ 4.1 Experiments on Different Datasets
In this section, to illustrate the universality of the proposed framework, we compare our proposed generalized framework, which combines standard first-order optimization algorithms SGD and second-order optimization algorithms oBFGS and oNAQ, with these algorithms on 12 different datasets. These datasets are diverse, including sparse, dense, imbalanced, high-dimensional, and low-dimensional. They were collected from the University of California, Irvine (UCI) Machine Learning repository [^1] https://archive.ics.uci.edu, Libsvm data [^2] https://www.csie.ntu.edu.tw/~cjlin/libsvmtools/datasets/. Details of the features are shown in Table 2. The second column indicates the source of the dataset, the third column indicates the quantity of the dataset, the fourth column indicates the dimension, and the fifth column indicates the quantity of data categories.
@@
The variance $\sigma$ in the AW function is set to 1 empirically. We conducted 100 iterative experiments, taking the total amount of iterations as the horizontal coordinate and the test accuracy as the vertical coordinate to depict the performance of each algorithm, as shown in Fig.7. Fig.7(a)-7(l) respectively plotted the experimental results of the 12 datasets in Table 2. No matter whether the proposed method in the initial result performs well or not, it tends to be stable after a certain number of iterations. We can conclude that the proposed approach effectively improves the performance of SGD, oBFGS and oNAQ on all 12 datasets.
@@
We additionally evaluate the efficacy of the proposed framework, AW-WSVM, by using the confusion matrix. To ensure the stability of the results, we iterated 100 times. The metrics of different algorithms obtained from the confusion matrix can be presented in Table 4-7, with the highest metrics bolded. Conclusions are easy to come by, and the proposed method outperforms SGD, oBFGS, and oNAQ on 12 datasets. The proposed method exhibits lower results on very few datasets. However, when considering the entirety of the recommended technique, it has successfully attained the intended outcome.
In order to intuitively illustrate the effectiveness of our proposed method in processing these 12 datasets, we plotted a histogram of G-mean values. We randomly sampled 20 experimental results for each dataset, and the height of each bar in the histogram represents the expectation of G-mean value over these 20 sampled results. As shown in Fig.8, we can easily observe that our method performs better on all datasets.
@@
Table 5: G-mean of six methods on different datasets.
Datasets oNAQ AW-WSVM+oNAQ oBFGS AW-WSVM+oBFGS SGD AW-WSVM+SGD
A7a 0.7726 0.7836 0.7854 0.7925 0.7998 0.8580
A8a 0.7838 0.7939 0.7882 0.7893 0.7969 0.8423
A9a 0.7869 0.7933 0.7830 0.7893 0.7710 0.8258
Mushroom 0.9804 0.9887 0.9804 0.9864 0.9083 0.9832
Yeast 0.7247 0.7515 0.7140 0.7534 0.5889 0.7590
Ijcnn1 0.6526 0.7863 0.7155 0.7906 0.5846 0.6726
W1a 0.8754 0.8949 0.8388 0.9495 0.6966 0.9009
W2a 0.8848 0.9228 0.8819 0.9241 0.6965 0.9189
W3a 0.9086 0.9483 0.9210 0.9654 0.6965 0.9375
W4a 0.9551 0.8997 0.8926 0.9825 0.6964 0.9501
W5a 0.9507 0.9356 0.8983 0.9397 0.6963 0.9693
W6a 0.9429 0.9445 0.9695 0.9262 0.6966 0.9787
Table 6: Recall of six methods on different datasets.
Datasets oNAQ AW-WSVM+oNAQ oBFGS AW-WSVM+oBFGS SGD AW-WSVM+SGD
A7a 0.8869 0.8900 0.8655 0.8846 0.7713 0.8769
A8a 0.8725 0.8800 0.8840 0.8859 0.7698 0.8846
A9a 0.8735 0.8825 0.8772 0.8814 0.7728 0.8745
Mushroom 0.9862 0.9908 0.9899 0.9863 0.8654 0.9832
Yeast 0.7333 0.7727 0.7142 0.7692 0.5000 0.7777
Ijcnn1 0.9056 0.9297 0.9058 0.9332 0.9049 0.9052
W1a 0.9769 0.9841 0.9734 0.9851 0.9706 0.9820
W2a 0.9756 0.9831 0.9769 0.9844 0.9703 0.9815
W3a 0.9748 0.9844 0.9736 0.9823 0.9702 0.9793
W4a 0.9755 0.9815 0.9752 0.9867 0.9702 0.9796
W5a 0.9728 0.9826 0.8983 0.9373 0.9699 0.9787
W6a 0.9762 0.9831 0.9732 0.9857 0.9707 0.9798
Table 7: F1-score of six methods on different datasets.
Datasets oNAQ AW-WSVM+oNAQ oBFGS AW-WSVM+oBFGS SGD AW-WSVM+SGD
A7a 0.8955 0.8996 0.8976 0.9025 0.8706 0.9038
A8a 0.8987 0.9030 0.9016 0.9022 0.8698 0.9046
A9a 0.9000 0.9033 0.8992 0.9018 0.8716 0.9048
Mushroom 0.9822 0.9897 0.9826 0.9874 0.9129 0.9850
Yeast 0.6241 0.6285 0.6277 0.6293 0.6004 0.6004
Ijcnn1 0.9500 0.9563 0.9502 0.9574 0.9501 0.9499
W1a 0.9876 0.9900 0.9864 0.9895 0.9706 0.9820
W2a 0.9872 0.9898 0.9878 0.9902 0.9849 0.9897
W3a 0.9870 0.9907 0.9865 0.9901 0.9848 0.9890
W4a 0.9874 0.9894 0.9868 0.9910 0.9848 0.9893
W5a 0.9861 0.9904 0.9856 0.9899 0.9847 0.9890
W6a 0.9876 0.9908 0.9863 0.9916 0.9851 0.9897
4.2 Statistical Comparison
$\chi_{F}^{2}=\frac{12N}{K(K+1)}[R_{j}^{2}-\frac{(K(K+1)^{2})}{4}]$ (15)
where $K$ is the number of algorithms involved in the comparison.
Table 8 shows the average ranking $R_{j}$ of the six classifiers in test accuracy. The p-value of Friedman’s test is $1.7122\times 10^{-9}$. The P-value is less than the threshold for significance level $\alpha=0.05$, indicating that the null assumption should be discarded. The above results show that there are great differences among all classification methods from a statistical point of view. We compared the g-mean of the six classifiers (see Table 9), with a p-value of $1.1560\times 10^{-5}$ for Friedman’s test, indicating significant differences among all classification methods. Table 10-11 record the Recall and F1-score of six classifiers. The Friedman’s test p-values for Recall and F1 are $1.4607\times 10^{-7}$ and $2.4246\times 10^{-8}$, respectively, which indicates significant differences between the methods.
@@ 4.2 Statistical Comparison
(a) A7a
(b) A8a
(c) A9a
(d) Mushroom
(e) Yeast
(f) Ijcnn1
(g) W1a
(h) W2a
(i) W3a
(j) W4a
(k) W5a
(l) W6a
@@
Fig.9(a)-9(d) respectively report the analysis results of six classifiers in terms of test accuracy, G-mean, Recall and F1-score. The test accuracy of the standard SGD, oBFGS, and oNAQ algorithms does not exhibit any substantial variation, but there is a notable disparity between the proposed approach in this research. Furthermore, our analysis indicates that the integration of all three algorithms with AW-WSVM leads to enhanced algorithm performance. This improvement is evident from the overall higher ranking of the combined algorithms across all datasets, showcasing varied progress. The results indicate a negligible disparity in the classification accuracy across standard SGD, oBFGS, and oNAQ. Each of the three algorithms was examined independently. The performance of SGD was enhanced by using the weight update approach suggested in this research (AW-WSVM+SGD). Notably, there were notable disparities between the two algorithms when compared to the original SGD. The same applies to second-order approaches such as BFGS and NAQ.
Table 9: Friedman ranking of these methods according to G-means on different datasets.
@@
A7a 6 5 4 3 2 1
A8a 6 3 5 4 2 1
A9a 4 2 5 3 6 1
Mushroom 4.5 1 4.5 2 5 3
Yeast 4 3 5 2 6 1
Ijcnn1 5 2 3 1 6 4
W1a 4 3 5 1 6 2
W2a 4 2 5 1 6 3
W3a 5 2 4 1 6 3
W4a 2 4 5 1 6 3
W5a 2 4 5 3 6 1
W6a 4 3 2 5 6 1
R 4.21 2.83 4.38 2.25 5.25 2.00
Table 10: Friedman ranking of these methods according to Recalls on different datasets.
Datasets oNAQ AW-WSVM+oNAQ oBFGS AW-WSVM+oBFGS SGD AW-WSVM+SGD
A7a 2 1 5 3 6 4
A8a 5 4 3 1 6 2
A9a 5 1 3 2 6 4
Mushroom 4 1 2 3 6 5
Yeast 4 2 5 3 6 1
Ijcnn1 4 2 3 1 6 5
W1a 4 2 5 1 6 3
W2a 5 2 4 1 6 3
W3a 4 1 5 2 6 3
W4a 4 2 5 1 6 3
W5a 3 1 6 5 4 2
W6a 4 2 5 1 6 3
R 4.00 1.75 4.25 2.00 5.83 3.16
Table 11: Friedman ranking of these methods according to F1-scores on different datasets.
Datasets oNAQ AW-WSVM+oNAQ oBFGS AW-WSVM+oBFGS SGD AW-WSVM+SGD
A7a 5 3 4 2 6 1
A8a 5 2 4 3 6 1
A9a 4 2 5 3 6 1
Mushroom 5 1 4 2 6 3
@@
Ijcnn1 5 2 3 1 4 6
W1a 3 1 4 2 6 5
W2a 5 2 4 1 6 3
W3a 4 1 5 2 6 3
W4a 4 2 5 1 6 3
W5a 4 1 5 2 6 3
W6a 4 2 5 1 6 3
R 4.33 1.75 4.25 1.75 5.79 3.12
(a) Classification accuracy
(b) Classification G-mean
(c) Classification Recall
(d) Classification F1-score
@@ 4.3 Experiments on Emotional Classification with Different Imbalance Ratio(IR)
In this section, we collect the dataset of emotion classification, which comes from gitte [^3] https://gitee.com/zhangdddong/toutiao-text-classfication-dataset. (see Tabel 12). The last column in the table indicates the imbalance ratio (IR). In daily life, different kinds of news data are always imbalanced, for example, entertainment news will account for more than science and technology news.
@@
In order to further illustrate the effectiveness of the proposed AW-WSVM framework, we process the same data and get G-means under different IRs[45]. Taking 102vs100 and 109vs100 as examples, we plot the results in Fig10. It is not difficult to find that our proposed method performs best in both low IR and high IR.
@@
5 Conclusion
Considering the massive data, we propose an Adaptive Weight function (AW function), through which the weights of samples can be dynamically updated. Our analysis demonstrates that the weights generated by the AW function obey a probability distribution in each class. We propose a new soft-margin weighted SVM and provide a more concise matrix expression, so that any stochastic optimization algorithm can iterate through the matrix, saving computational space. Since minority is extremely sensitive to noise, we propose an effective noise filtering method based on nearest neighbors, which improves the robustness of the model. Subsequently, we can allocate greater weights to samples that are in close proximity to the decision hyperplane, while assigning lower weights to the other samples. Our allocation technique is appropriate since the support vectors, which are placed near the decision hyperplane, correspond with the concept that samples close to the decision plane have higher influence.
@@ 5 Conclusion
The proposed framework, AW-WSVM, was evaluated on a total of 12 different datasets sourced from UCI repositories and LIBSVM repositories. The experiments conducted confirm that the proposed framework AW-WSVM surpasses the conventional first-order approaches SGD, second-order methods oBFGS, and oNAQ. Furthermore, it exhibited superior performance in terms of classification accuracy, g-mean, recall and F1-score. In addition, we also conducted statistical experiments on these six classifiers. Friedman tests and Nemenyi post hoc tests demonstrate that the standard algorithms SGD, oBFGS, and oNAQ all perform better within our proposed framework. We not only do experiments on standard datasets, but also apply the framework to emotion classification datasets. From the G-means and other indicators, the framework proposed in this paper can effectively improve the performance of standard optimization algorithms, no matter whether the imbalance index is high or low. Moreover, the core of the proposed generalized framework, AW-WSVM, is to look for data samples near the classified hyperplane. Consequently, it may be effectively integrated with any soft-margin SVM solution techniques.